# HDC: Hierarchical Distillation for Multi-level Noisy Consistency in Semi-Supervised Fetal Ultrasound Segmentation

### v1

@@ Abstract
Transvaginal ultrasound is a critical imaging modality for evaluating cervical anatomy and detecting physiological changes. However, accurate segmentation of cervical structures remains challenging due to low contrast, shadow artifacts, and fuzzy boundaries. While convolutional neural networks (CNNs) have shown promising results in medical image segmentation, their performance is often limited by the need for large-scale annotated datasets—an impractical requirement in clinical ultrasound imaging. Semi-supervised learning (SSL) offers a compelling solution by leveraging unlabeled data, but existing teacher-student frameworks often suffer from confirmation bias and high computational costs. We propose HDC, a novel semi-supervised segmentation framework that integrates Hierarchical Distillation and Consistency learning within a multi-level noise mean-teacher framework. Unlike conventional approaches that rely solely on pseudo-labeling, we introduce a hierarchical distillation mechanism that guides feature-level learning via two novel objectives: (1) Correlation Guidance Loss to align feature representations between the teacher and main student branch, and (2) Mutual Information Loss to stabilize representations between the main and noisy student branches. Our framework reduces model complexity while improving generalization. Extensive experiments on two fetal ultrasound datasets, FUGC and PSFH, demonstrate that our method achieves competitive performance with significantly lower computational overhead than existing multi-teacher models.
@@ 1 Introduction
In contrast, we propose a hierarchical distillation mechanism that enriches the learning signal and improves feature alignment. This is achieved through two novel loss functions: (1) Correlation Guidance Loss, which encourages consistency between the feature representations of the teacher and the main student branch, and (2) Mutual Information Loss, which aligns the representations between the main student and a noisy student branch to enhance intra-batch stability. These mechanisms are designed to bridge the feature distribution gap between teacher and student networks while stabilizing student representations in the presence of noise. HDC is evaluated on two challenging fetal ultrasound datasets, FUGC and PSFH [1, 21], where it achieves state-of-the-art segmentation performance with significantly reduced computational overhead compared to existing methods under various evaluation metrics.
@@ 2 Related Works
Knowledge Distillation (KD). Originally introduced by Hinton et al. [10], this technique enables the transfer of knowledge from a large, complex teacher model to a smaller student model, enabling efficient deployment while maintaining performance. KD plays important roles in opening new approaches for many fields such as multi-modal tasks[15, 30, 41]. Originally proposed as a method where the student mimics the teacher’s soft predictions, KD has since evolved into multiple paradigms, each capturing different aspects of knowledge transfer. KD methods are commonly categorized into logits-based [10, 14], feature-based [3, 4, 9], and relation-based approaches [12, 24]. Response-based distillation aligns the student’s predictions with the teacher’s soft outputs, often using Kullback-Leibler [42] divergence. Beyond classical approaches, recent works have explored alternative formulations grounded in self-supervision and information theory, aiming for a more structured and effective knowledge transfer. Some methods adopt orthogonal projection techniques [25] to distill knowledge more efficiently, while others focus on activation gradient alignment to better capture the loss landscape [38].
@@ 3.3.2 Mutual Information Loss
To further enhance knowledge transfer, we move beyond feature-wise alignment and incorporate intra-batch consistency through a mutual information loss. While the correlation loss ensures that the student’s main decoder is well-aligned with the teacher, it does not explicitly enforce consistency between the student’s two decoders. To address this, we establish a hierarchical distillation scheme, where knowledge is not only transferred from the teacher to the main decoder but also propagated from the main decoder to the noisy one. This structured flow of information ensures that both decoders develop robust representations while simultaneously allowing the shared encoder to learn complementary features from both branches.
@@ 4.2 Implementation Details
For the PSFH Dataset, we used the same setting as DSTCT [13]. Using the Stochastic Gradient Descent (SGD) optimizer set up with a momentum of 0.9 and a weight decay of 0.0001, the network training protocol involved 30,000 iterations. To accommodate semi-supervised learning paradigms, we used a batch size 16, with an equal distribution of eight tagged and eight unlabeled photos. A clustering-based learning rate technique was used to modify the training’s initial learning rate of 0.01 dynamically. Data noise perturbation in the [-0.2, 0.2] range was added to improve the model’s resilience to input variability
@@ Acknowledgement
We acknowledge the support from University of Information Technology (UIT), and AI VIETNAM Lab for providing the GPU for the numerical calculations.

### v2

@@ Abstract
Transvaginal ultrasound is a critical imaging modality for evaluating cervical anatomy and detecting physiological changes. However, accurate segmentation of cervical structures remains challenging due to low contrast, shadow artifacts, and indistinct boundaries. While convolutional neural networks (CNNs) have demonstrated efficacy in medical image segmentation, their reliance on large-scale annotated datasets presents a significant limitation in clinical ultrasound imaging. Semi-supervised learning (SSL) offers a potential solution by utilizing unlabeled data, yet existing teacher-student frameworks often encounter confirmation bias and high computational costs. In this paper, a novel semi-supervised segmentation framework, called HDC, is proposed incorporating adaptive consistency learning with a single-teacher architecture. The framework introduces a hierarchical distillation mechanism with two objectives: Correlation Guidance Loss for aligning feature representations and Mutual Information Loss for stabilizing noisy student learning. The proposed approach reduces model complexity while enhancing generalization. Experiments on fetal ultrasound datasets, FUGC and PSFH, demonstrate competitive performance with reduced computational overhead compared to multi-teacher models.
@@ 1 Introduction
In contrast, we propose a hierarchical distillation mechanism that enriches the learning signal and improves feature alignment, inspired by theoretical insights from prior works on knowledge distillation [49, 25]. This is achieved through two novel loss functions: (1) Correlation Guidance Loss, which encourages consistency between the feature representations of the teacher and the main student branch, and (2) Mutual Information Loss, which aligns the representations between the main student and a noisy student branch to enhance intra-batch stability. These mechanisms are designed to bridge the feature distribution gap between teacher and student networks while stabilizing student representations in the presence of noise. HDC is evaluated on two challenging fetal ultrasound datasets, FUGC and PSFH [1, 21], where it achieves state-of-the-art segmentation performance with significantly reduced computational overhead compared to existing methods under various evaluation metrics.
@@ 2 Related Works
Knowledge Distillation (KD). Originally introduced by Hinton et al. [10], this technique enables the transfer of knowledge from a large, complex teacher model to a smaller student model, enabling efficient deployment while maintaining performance. KD plays important roles in opening new approaches for many fields such as multi-modal tasks[15, 31, 42]. Originally proposed as a method where the student mimics the teacher’s soft predictions, KD has since evolved into multiple paradigms, each capturing different aspects of knowledge transfer. KD methods are commonly categorized into logits-based [10, 14], feature-based [3, 4, 9], and relation-based approaches [12, 24]. Response-based distillation aligns the student’s predictions with the teacher’s soft outputs, often using Kullback-Leibler [43] divergence. Beyond classical approaches, recent works have explored alternative formulations grounded in self-supervision and information theory [25], aiming for a more structured and effective knowledge transfer. Some methods adopt orthogonal projection techniques [26] to distill knowledge more efficiently, while others focus on activation gradient alignment to better capture the loss landscape [39].
@@ 3.3.2 Mutual Information Loss
To further enhance knowledge transfer, we move beyond feature-wise alignment and incorporate intra-batch consistency through a mutual information loss, originally proposed in [25]. While the correlation loss ensures that the student’s main decoder is well-aligned with the teacher, it does not explicitly enforce consistency between the student’s two decoders. To address this, we establish a hierarchical distillation scheme, where knowledge is not only transferred from the teacher to the main decoder but also propagated from the main decoder to the noisy one. This structured flow of information ensures that both decoders develop robust representations while simultaneously allowing the shared encoder to learn complementary features from both branches.
@@ 4.2 Implementation Details
For the PSFH Dataset, the same setting as DSTCT [13] was used. Using the Stochastic Gradient Descent (SGD) optimizer set up with a momentum of 0.9 and a weight decay of 0.0001, the network training protocol involved 30,000 iterations. To accommodate semi-supervised learning paradigms, we used a batch size 16, with an equal distribution of eight tagged and eight unlabeled photos. A clustering-based learning rate technique was used to modify the training’s initial learning rate of 0.01 dynamically. Data noise perturbation in the [-0.2, 0.2] range was added to improve the model’s resilience to input variability
@@ Acknowledgement
This research is supported from VNUHCM - University of Information Technology (UIT). AI VIETNAM Lab supports for providing the GPU to numerical calculate.